OpenStudy (anonymous):

Every month, a girl gets allowance. Assume last year she had no money, and kept it up to now. Then she spends 1/2 of her money on clothes, then 1/3 of the remaining money on games, and then 1/4 of the remaining money on toys. After she bought all of that, she had $7777 left. Assuming she only gets money by allowance, how much money does she earn every month?

OpenStudy (anonymous):

plz answer

OpenStudy (anonymous):

the question is what portion of her money did she spend?

OpenStudy (anonymous):

ok so how much does she earn

OpenStudy (anonymous):

okay actually that was wrong she spends \(\frac{1}{2}\) of her money on clothes, leaving \(\frac{1}{2}\) of her money left then she spend \(\frac{1}{3}\) of that on games, and \(\frac{1}{2}\times \frac{1}{3}=\frac{1}{6}\) so she has \(\frac{1}{2}-\frac{1}{6}=\frac{1}{3}\) of her money left

OpenStudy (anonymous):

then she spends \(\frac{1}{4}\) of that on toys and \(\frac{1}{4}\times \frac{1}{3}=\frac{1}{12}\)

OpenStudy (anonymous):

so its $2222 right?

OpenStudy (anonymous):

leaving her with \[\frac{1}{3}-\frac{1}{12}=\frac{1}{4}\] of her original mount of $

OpenStudy (anonymous):

now the question is 7777 is one fourth of what number?

OpenStudy (anonymous):

which is another way of asking, what is \(4\times 7777\)

OpenStudy (anonymous):

2222?

OpenStudy (anonymous):

31108

OpenStudy (anonymous):

no, the number has to be larger than 7777

OpenStudy (anonymous):

that is what i get, yes